OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

J=T+10 and 2J=T-3

OpenStudy (anonymous):

now 3 yrs ago coach 10+x 7+x tim x x-3 7+x=2(x-3) 7+x=2x-6 x=13